Find the probability that a dart will (a)hit in the baseball field (b) hit in one of the 4 smaller squares (c) will miss all five target squares, on a square dart board that has 60cm sides with a square baseball field (inside the square dart board)40cm on it's sides, and 4 smaller squares (inside square dart board, but outside of baseball field) each 10cm on a side as the targets.
Answer by stanbon(29457)

Find the probability that a dart will
(a)hit in the baseball field = 1600/3600 = 4/9
(b) hit in one of the 4 smaller squares = 400/3600 = 1/9
(c) will miss all five target squares: 4/9
-------------------------------------------
on a square dart board that has 60cm sides: area = 3600 sq cm.
-----------------------------------
square field (inside the dart board)40cm on it's sides: 40^2 = 1600 sq cm.
----------------------,
and 4 smaller squares (inside square dart board, but outside of baseball field) each 10cm on a side as the targets. 4*10^2 = 400 sq cm.
------------------------
Cheers,
Stan H.